OpenStudy (anonymous):

5x+3y=15 find the x intercept and the y intercept of the graph of each equation

OpenStudy (striker1298):

jose each?

OpenStudy (striker1298):

I see only one equation

OpenStudy (striker1298):

the y-intercept is (0,5) and the x-intercept is (3,0)

OpenStudy (anonymous):

2x-6y=12 3x-4y-10=0

OpenStudy (anonymous):

2x+5y-10=0 y=x y=4x-2 this ones too :)

OpenStudy (striker1298):

joselyni just go on https://www.desmos.com/calculator it will graph it for you and give you the y-intercept and the x-intercept gotta go good luck